# Statistical Mechanics of the Sub-Optimal Transport

### v1

@@ 1 Introduction
These methods have been also successfully applied to a broad class of combinatorial optimization problems defined on graphs. Among the most studied are matching problems, where the goal is to pair elements across two sets at minimum cost. Mézard and Parisi applied the replica method to compute the expected cost of optimal matchings in both the low-temperature limit and the thermodynamic limit  [13, 14], while Aldous found a rigorous derivation in the same regime [1]. Related problems include stable marriage and the roommate problem, where preferences impose additional stability constraints [6], and the traveling salesman problem, which seeks an optimal cycle through a set of cities [12, 19]. More recently, [2] reformulated matching within a grand-canonical framework, recovering the classical Mézard-Parisi results while sharing some mathematical tools with the model described in this work.
@@
A common feature unites these analyses: they target the zero-temperature limit, where the system collapses onto a single optimal configuration amenable to exact treatment. Real-world systems, however, rarely achieve perfect optimization. Configurations that are sub-optimal yet structured, neither fully random nor fully optimized, emerges frequently. The analytical machinery developed for zero-temperature problems does not directly apply to this intermediate regime, because taking this limit before the thermodynamic limit eliminates precisely the sub-optimality of interest. A statistical mechanics framework capable of describing systems at finite temperature, where entropy and cost minimization genuinely compete, remains a gap in the literature.
The Sub-Optimal Transport (SOT) model introduced in [3] provides a natural setting to address this challenge. The model defines an ensemble of weighted bipartite graphs in which a tuning parameter $\beta$ interpolates continuously between entropy-dominated (dense, random) and cost-dominated (sparse, structured) configurations. Numerical simulations reveal a well-defined crossover between these regimes, yet no analytical theory has captured this transition within a statistical mechanics framework. Providing such a theory would clarify how large-scale structural organization emerges from the competition between randomness and optimization.
@@ 2 Sub-OT model and the relation between stat-mech and maximum entropy
These constraints define an ensemble of random weighted bipartite graphs, where each $W_{i\alpha}$ represents mass on the link connecting node $i$ in one layer to node alphaalpha in the other. The cost matrix $C_{i\alpha}$ enters not as a constraint but through Boltzmann-like weighting, yielding the following probability distribution over mass configurations:
@@
provides a natural point of comparison emerging from the canonical structure of Eq.(3). The statistical mechanics approach evaluates the partition function $Z$ by integrating over all variables — $W$, $\lambda$, and $\mu$ — thereby accounting for all configurations consistent with constraints. In contrast, the information-based approach integrates only over $W$ while fixing $\lambda$ and $\mu$ at their most probable values (the maximum of the log-likelihood). This constitutes an approximation that converges to the full $Z$ only when $\mathcal{F}$ in Eq.(4) exhibits a dominant peak around the most probable configuration.
@@ 3.1 Total mass conservation
Here $\bar{\omega}=\frac{\bar{W}}{N_{\gamma}}$ is the average mass per edge and the free energy density is:
@@
Second, the free energy density $\mathrm{F}$ remains analytic and smooth at all orders of $\beta$. The system therefore does not undergo a true thermodynamic phase transition but rather exhibits a continuous crossover between regimes, consistent with the absence of any singularities in $\mathrm{F}$.
@@
For $N_{\gamma}\gg 1$ the integrand becomes sharply peaked around its maximum. The integral is therefore dominated by the value of $\lambda$ that maximizes the exponent—the saddle point $\lambda^{*}$. Taking the derivative of the argument of the exponential with respect to $\lambda$ and setting it to zero yields:
@@
This observable, shown in panel (b) of Figure 1, shows two distinct regimes: In the small $\beta$ (high-noise) regime entropy dominates and costs decorrelate from weights. The system approaches a uniform distribution of weights, where each edge carries approximately $\bar{\omega}\sim W_{i\alpha}/N_{\gamma}$ independently of its cost $C_{i\alpha}$. Since costs are drawn uniformly from $[0,1]$, the average energy factorizes as $\langle cW\rangle\to\langle c\rangle\,\bar{\omega}=\bar{\omega}/2$ where $\langle c\rangle=1/2$ is the mean of the cost distribution. As $\beta$ increases mass migrates from high-cost edges $(C_{i\alpha}\sim 1)$ to low-cost edges $(C_{i\alpha}\to 0)$. The average energy vanishes asymptotically, $u\to 0$, because the probability distribution allocates mass exclusively to the subset of edges with minimal cost.
@@ 3.2 Full model
The full model requires node-wise conservation at both layers. Beyond the total mass constraint in Eq.(16), the model imposes $N$ constraints ensuring that each source node $i$ emits exactly $s_{i}$ units of mass and $M$ constraints ensuring that each target node $\alpha$ receives exactly $\sigma_{\alpha}$. Together, these three constraint types—one global and $N+M$ local—fully specify the ensemble of admissible mass distributions.
@@
Before proceeding further, we examine numerical solutions of the SOT model to understand the behavior of local fluctuations. The partition function in its current form depends on $N+M$ fluctuation variables whose distributions are not known a priori. To guide our analytical treatment, we numerically solve the SOT model by finding the Lagrange multipliers $\{\lambda_{i},\mu_{\alpha}\}$ that satisfy the saddle-point conditions for various system sizes and values of $\beta$. Figure 2 presents the results across different regimes.
@@
The variance exhibits more intricate behavior: it increases with $\beta$ while decreasing with system size (panel d). The inset of panel (d) reveals the critical non-commutativity of limits through a vertical slice at high $\beta$: the variance follows a power-law decay with exponent $\sim-1$ as system size grows. This demonstrates that the order of limits matters crucially. If the large-$\beta$ limit is taken first ($\beta\to\infty$ before $N,M\to\infty$), the variance grows indefinitely without being compensated by system-size effects, and fluctuations remain large. Conversely, if the thermodynamic limit is taken first, the variance vanishes and the Gaussian approximation becomes exact. In what follows, we always take the thermodynamic limit first and only afterwards consider the large-$\beta$ regime.
@@
where $\Psi(l,m,s,\sigma,\beta C)$ collects all the fluctuating terms involving Lagrange multipliers. As shown in B, $\Psi$ is sub-extensive in $(N,M)$ and can therefore be neglected from the elements affecting the partition function in the thermodynamics limit. The partition function reduces to the one in Eq.(8):
@@
This demonstrates that, in the mean-field framework, the partition function of the SOT model with all the strengths constraints is dominated by the partition function of the model with only the global constraint, discussed in Section 3.1. The full complexity of local fluctuations becomes irrelevant in the thermodynamic limit, and the essential physics is captured by the simpler global-constraint model.
@@ 3.3 Generalization to other homogeneous cost distributions
For costs drawn from a uniform distribution, the average energy follows Eq.(14). Panel (a) of Figure 3 shows excellent agreement between numerical results and the theoretical prediction across all system sizes for $\bar{\omega}=1$. This agreement holds also for all constraint values $\bar{\omega}$ (not shown).
The generalization to non-uniform cost distributions requires only numerical evaluation of the saddle point while preserving the analytical structure. For a general cost distribution $\rho(c)$, we can start from Eq.(8) and (9) since the partition function retains the form:
@@
$0=\frac{\partial}{\partial\lambda}[\bar{\omega}\lambda-\Phi(\lambda,\beta)]\quad\Rightarrow\quad\bar{\omega}=\int\rho(c)\frac{1}{\beta c+\lambda}dc$ (22)
@@
Panel (b) of Figure 3 shows the comparison for costs drawn from a Beta distribution $\rho(c)\sim c^{\eta}$ with support $c\in[0,1]$. In the SOT model, the exponent $\eta$ depends on system size through $\eta=1/\log_{10}(L)$, which explains the discrepancy observed in the dense regime where finite-size effects remain significant. As expected, the agreement becomes excellent in the large-$\beta$ limit while finite-size effects become apparent at lower $\beta$ values (see inset).
@@ 4 Weight distribution in the ensemble and analysis in the large-$\beta$ regime
The SOT model provides a framework to connect observable weight patterns to underlying optimization constraints. Notice that the cost structure driving these patterns may be unknown or uncertain. However, the weight distribution $\rho_{W}(w)$ and the cost distribution $\rho_{C}(c)$ are intrinsically linked: properties of the cost distribution manifest directly as observable features in the weight statistics. In this section we demonstrate that, for uniformly distributed costs, the weight distribution follows a power law $\rho_{W}(w)\sim w^{-2}$ in the large-$\beta$ regime.
@@
At large $\beta$, the optimization selects a small subset of edges with minimal costs to carry most of the mass, while weights on the remaining unselected edges become vanishingly small. For these unselected edges, the weight is so small that it is plausibly independent of the specific cost value—what matters is only that the edge was not selected. Crucially, the number of selected edges scales linearly with system size $N$ (forming a spanning tree), while the number of unselected edges grows quadratically as $N^{2}$. Since the empirical weight distribution is dominated by the much larger population of unselected edges, it is reasonable to assume that the Lagrange multipliers $\theta_{i\alpha}$ are uncorrelated with the corresponding costs $C_{i\alpha}$ in this regime. Under this assumption, the joint distribution factorizes $\rho(c,\theta)\approx\rho_{C}(c)\,\rho_{\theta}(\theta)$. The distribution $\rho_{\theta}(\theta)$ is the Gaussian established in Figure 2 panel (b). This factorization allows us to compute the weight distribution $\rho_{W}(w)$ by convolving these independent distributions:

### v2

@@ 1 Introduction
These methods have also been successfully applied to a broad class of combinatorial optimization problems defined on graphs. Among the most studied are matching problems, where the goal is to pair elements across two sets at minimum cost. Mézard and Parisi applied the replica method to compute the expected cost of optimal matchings in both the low-temperature limit and the thermodynamic limit  [13, 14], while Aldous found a rigorous derivation in the same regime [1]. Related problems include stable marriage and the roommate problem, where preferences impose additional stability constraints [6], and the traveling salesman problem, which seeks an optimal cycle through a set of cities [12, 19]. More recently, [2] reformulated matching within a grand-canonical framework, recovering the classical Mézard-Parisi results while sharing some mathematical tools with the model described in this work.
@@
A common feature unites these analyses: they target the zero-temperature limit, where the system collapses onto a single optimal configuration amenable to exact treatment. Real-world systems, however, rarely achieve perfect optimization. Configurations that are sub-optimal yet structured, neither fully random nor fully optimized, emerges frequently. The analytical machinery developed for zero-temperature problems does not directly apply to this intermediate regime, because taking this limit before the thermodynamic limit eliminates precisely the sub-optimality of interest: the thermodynamic limit and the zero-temperature limit do not commute. A statistical mechanics framework capable of describing systems at finite temperature, where entropy and cost minimization genuinely compete, remains a gap in the literature.
The Sub-Optimal Transport (SOT) model, introduced in [3], provides a natural setting to address this challenge. The model defines an ensemble of weighted bipartite graphs in which a tuning parameter $\beta$ interpolates continuously between entropy-dominated (dense, random) and cost-dominated (sparse, structured) configurations. Numerical simulations reveal a well-defined crossover between these regimes, yet no analytical theory has captured this transition within a statistical mechanics framework. Providing such a theory would clarify how large-scale structural organization emerges from the competition between randomness and optimization.
@@ 2 Sub-OT model and the relation between stat-mech and maximum entropy
These constraints define an ensemble of random weighted bipartite graphs, where each $W_{i\alpha}$ represents mass on the link connecting node $i$ in one layer to node $\alpha$ in the other. The cost matrix $C_{i\alpha}$ enters not as a constraint but through Boltzmann-like weighting, yielding the following probability distribution over mass configurations:
@@
provides a natural point of comparison emerging from the canonical structure of Eq.(3). The statistical mechanics approach evaluates the partition function $Z$ by integrating over all variables ($W$, $\lambda$, and $\mu$) thereby accounting for all configurations consistent with constraints. In contrast, the information-based approach integrates only over $W$ while fixing $\lambda$ and $\mu$ at their most probable values (the maximum of the log-likelihood). This constitutes an approximation that converges to the full $Z$ only when $\mathcal{F}$ in Eq.(4) exhibits a dominant peak around the most probable configuration.
@@ 3.1 Total mass conservation
Here $\bar{\omega}=\frac{\bar{W}}{N_{\gamma}}$ is the average mass per edge and the free energy is:
@@
Second, the free energy $\mathrm{F}$ remains analytic and smooth at all orders of $\beta$. The system therefore does not undergo a true thermodynamic phase transition but rather exhibits a continuous crossover between regimes, consistent with the absence of any singularities in $\mathrm{F}$.
@@
For $N_{\gamma}\gg 1$, the integrand becomes sharply peaked around its maximum. The integral is therefore dominated by the value of $\lambda$ that maximizes the exponent, i.e. the saddle point $\lambda^{*}$. Taking the derivative of the argument of the exponential with respect to $\lambda$ and setting it to zero yields:
@@
This observable, shown in panel (b) of Figure 1, shows two distinct regimes: In the small $\beta$ (high-noise) regime, entropy dominates and costs decorrelate from weights. The system approaches a uniform distribution of weights, where each edge carries approximately $\bar{\omega}\sim W_{i\alpha}/N_{\gamma}$ independently of its cost $C_{i\alpha}$. Since costs are drawn uniformly from $[0,1]$ and are independent, the average energy factorizes as $\langle cW\rangle\to\langle c\rangle\,\bar{\omega}=\bar{\omega}/2$ where $\langle c\rangle=1/2$ is the mean of the cost distribution. As $\beta$ increases mass migrates from high-cost edges $(C_{i\alpha}\sim 1)$ to low-cost edges $(C_{i\alpha}\to 0)$. The average energy vanishes asymptotically, $u\to 0$, because the probability distribution allocates mass exclusively to the subset of edges with minimal cost.
@@ 3.2 Full model
The full model requires node-wise conservation at both layers. Beyond the total mass constraint in Eq.(16), the model imposes $N$ constraints ensuring that each source node $i$ emits exactly $s_{i}$ units of mass and $M$ constraints ensuring that each target node $\alpha$ receives exactly $\sigma_{\alpha}$. Together, these three constraint types (one global and $N+M$ local) fully specify the ensemble of admissible mass distributions.
@@
Before proceeding further, we examine numerical solutions of the SOT model to understand the behavior of local fluctuations. The partition function in its current form depends on $N+M$ fields variables whose distributions are not known a priori. To guide our analytical treatment, we numerically solve the SOT model by finding the Lagrange multipliers $\{\lambda_{i},\mu_{\alpha}\}$ that satisfy the saddle-point conditions for various system sizes and values of $\beta$. Figure 2 presents the results across different regimes.
@@
The variance exhibits more intricate behavior: it increases with $\beta$ while decreasing with system size (panel d). The inset of panel (d) reveals the critical non-commutativity of limits through a vertical slice at high $\beta$: the variance follows a power-law decay with exponent $\sim-1$ as system size grows. This demonstrates that the order of limits matters crucially. If the large-$\beta$ limit is taken first ($\beta\to\infty$ before $N,M\to\infty$), the variance grows indefinitely without being compensated by system-size effects, and fluctuations remain large. Conversely, if the thermodynamic limit is taken first, the variance vanishes, and the Gaussian approximation becomes exact. In what follows, we always take the thermodynamic limit first since we are interested in the finite $\beta$ regime.
@@
where $\Psi(l,m,s,\sigma,\beta C)$ collects all the fluctuating terms involving Lagrange multipliers. As discussed in B, $\Psi$ is sub-extensive in $(N\times M)$ and can therefore be neglected from the elements affecting the partition function in the thermodynamics limit. The partition function reduces to the one in Eq.(8):
@@
This demonstrates that, in the mean-field framework, the partition function of the SOT model with all the strength constraints is dominated by the partition function of the model with only the global constraint, discussed in Section 3.1. The full complexity of local fluctuations becomes irrelevant in the thermodynamic limit, and the essential physics is captured by the simpler global-constraint model.
For costs drawn from a uniform distribution, the average energy converges to the prediction of Eq.(14). Panel (a) of Figure 3 shows excellent agreement between numerical results and the theoretical prediction across all system sizes, in both the large and low $\beta$ regimes.
@@ 3.3 Generalization to other homogeneous cost distributions
The generalization to non-uniform but homogeneous cost distributions is possible using the ensemble equivalence, allowing the computation of the free energy through saddle point. For a general cost distribution $\rho(c)$, we can start from Eq.(8) and (9) since the partition function retains the form 333here we consider a complex rotation $\lambda\to\imath\lambda$.:
@@
$0=\frac{\partial}{\partial\lambda}[\bar{\omega}\lambda-\Phi(\lambda,\beta)]\Big|_{\lambda^{*}}\quad\Rightarrow\quad\bar{\omega}=\int\rho(c)\frac{1}{\beta c+\lambda^{*}}dc$ (22)
@@
Panel (b) of Figure 3 shows the comparison for costs drawn from a Beta distribution $\rho(c)\sim c^{\eta}$ with support $c\in[0,1]$. This behavior is representative of the broader class of cost laws with vanishing density at $c\to 0$, and might be therefore qualitatively distinct from the uniform case, which instead typifies distributions with finite $\rho(0)$ (see [1] for a short discussion). In the simulations of the SOT model considered in [3], the exponent $\eta$ depends on system size through $\eta=1/\log_{10}(L)$, which explains the discrepancy observed in the dense regime where finite-size effects remain significant. As expected, the agreement becomes excellent in the large-$\beta$ regime while finite-size effects become apparent at lower $\beta$ values, but vanish as the size increases (see inset).
@@ 4 Weight distribution in the ensemble and analysis in the large-$\beta$ regime
The SOT model provides a framework to connect observable weight patterns to underlying optimization constraints. Notice that the cost structure driving these patterns may be unknown or uncertain. However, the weight distribution $\rho_{W}(w)$ and the cost distribution $\rho_{C}(c)$ are intrinsically linked: properties of the cost distribution manifest directly as observable features in the weight statistics. In this section we demonstrate that, for short-tails distributed costs, the weight distribution follows a power law $\rho_{W}(w)\sim w^{-2}$ in the large-$\beta$ regime.
@@
At large $\beta$, the optimization selects a small subset of edges with minimal costs to carry most of the mass, while weights on the remaining unselected edges become vanishingly small. For these unselected edges, the weight is so small that it is plausibly independent of the specific cost value, what matters is only that the edge was not selected. Crucially, the number of selected edges scales linearly with system size $N+M$ (forming a spanning tree), while the number of unselected edges grows quadratically as $N\times M$. Since the empirical weight distribution is dominated by the much larger population of unselected edges, it is reasonable to assume that the Lagrange multipliers $\theta_{i\alpha}$ are uncorrelated with the corresponding costs $C_{i\alpha}$ in this regime. Under this assumption, the joint distribution factorizes $\rho(c,\theta)\approx\rho_{C}(c)\,\rho_{\theta}(\theta)$. The distribution $\rho_{\theta}(\theta)$ is the Gaussian established in Figure 2 panel (b). This factorization allows us to compute the weight distribution $\rho_{W}(w)$ by convolving these independent distributions:
@@ 5 Conclusion
Acknowledgemens
D.M., A.P. and R.P. acknowledge the financial support under the National Recovery and Resilience Plan (NRRP), Mission 4, Component 2, Investment 1.1, Call for tender No. 104 published on 2.2.2022 by the Italian Ministry of University and Research (MUR), funded by the European Union – NextGenerationEU– Project Title "WECARE – WEaving Complexity And the gReen Economy" – CUP 20223W2JKJ by the Italian Ministry of University and Research (MUR).